From GlueXWiki
Jump to: navigation, search

October 11, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Status
    • First package assembly
  2. Cathode corrosion test results (Vlad)
  3. Engineering: temperature control unit(Bill)
  4. Electronics (Nick, Fernando)
  5. Third package tests FDC E-log (Beni)
  6. TDR, pages 132-146 (Lubomir)
  7. Other


Participants: Fernando, Eugene, Bill, Dave, Nick, Vlad, Simon, Beni, and Lubomir. This was the longest "short" meeting (Dave).


- Dave: Fourth cell of the first package was installed yesterday, today we put a cathode from the fifth cell, tomorrow we plan to finish also the last cell. We are using the six spare cathodes made at the end of the production. The last cathode (#58) to finish the package needs the cards glued on it. Wire frame #29 is ready, except wire deadening. Wire frame #30 is on the stringing table, we will have Casey and Mike later in the month to string it.

- Now we are at most four people at Blue Crab: Dave, Anatoly, Vlad, and Lubomir. Moving the cathodes with the vacuum fixture require four people.

- Eugene: what needs to be ordered till the end of the FDC production. Dave: rigid-flex population (connectors ordered), grounding clamps, plus consumables. All the rest has been ordered.

Cathode corrosion

- Vlad showed results of the tests in the vacuum chamber at high temperature (~100degC) where all the samples (cathode traces 0.5mm wide) have wires soldered and he measures their resistance continuously without opening the chamber. Using gas, the EPDM sample showed >1MOhm resistance after 8-10 days, which translated to room temperature is ~100 days or more depending on the activation energy. No changes in the Viton, and Viton+Apiezon resistance after 15 days. New tests started with air at 100degC and 100% humidity. No effects yet after 3.5 days.

- Eugene: humidity doesn't seem to be such important factor; if that's the case the gas itself can contribute significantly to the corrosion? But then one has to explain the Viton/ EPDM difference, and why the corrosion happens at the O-ring. Eugene proposed to do also experiment in air but at low humidity. In the moment the priority of the studies is to investigate the Viton+Apiezon samples at worse possible conditions and then translate the results to the hall conditions.

- Dave found that the CO2 used by the gas company to make our mixture at Blue Crab is grade 3 (Argon is grade 5), meaning contaminations can be at 0.1% level and are not defined. Dave will talk to the company and fix this.


- Bill proposes an off-shell Temperature Control Unit (document linked above) to be used in the FDC cooling system. The unit is designed for Fluorinert and is ready for use. It will need probably some extra stuff, like flowmeters to detect leakages, RTDs to monitor the temperature, but in general it will work. Bill had some concerns also about the type of the pump used, that can create high pressure, so one has to find a way to control it, as well. The price is similar to the designed unit, that was supposed to be made at home. After some discussions Eugene agreed the off-shell Unit to be purchased: it will save time.

- What will happen if the Unit fails? Fernando: the voltage regulator on the pre-amp card has over-current and over-temperature protection, i.e. it will shut off the power to the card. In addition the control system should shut off the LV if the Temperature Control Unit doesn't work.


- Nick checked the board that was noisy at Blue Crab, couldn't see the noise, just one channel had on offset; will test it again at Blue Crab.

- Fernando initiated discussion about the plans for the FDC and CDC testing (at the end of the meeting). According to the contract we are suppose to have first article (10 units initially, but later not specified in the contract) fADC125 in January, then 2 weeks to approve, then the full production will go till end of March/April.

- The plan for the FDC is to do partial testing of all the packages at 126 connecting all the wires to TDCs, but using only two fADC125 to test the cathodes. The full test of the packages (need 36 fADCs) is scheduled to start in April and, assuming one month per package, will finish by August. Then the packages will be installed on the cart at Blue Crab and moved to the Hall in the fall.

- Fernando: what if there's a delay with the fADCs? We decided that we can test the first package(s) with half of the fADCs (18) and use 36 fADCS when we have them, for the rest of the packages. At least one has to be tested with all the fADCs. For that Fernando will talk to the company to have 20 fADCs as first article.

- fADC firmware: should not rely on that for the tests, except we should be able to change the addresses of the modules, now hard-coded. We should make contract with Gerard to develop the firmware.

Third package tests

- Temperature at 126 is very high since Monday, today it was 86F! Facilities work on that issue.

- Beni showed latest results (entry 89 and 90 in FDC Elog). Trigger window didn't match exactly the TDC signals, therefore the very beginning of the drift spectra was missing. Now the efficiency is ~100%. Another issue is the hit multiplicity in the TDCs. It's about 3 changing with the HV (entry 90). It's much lower for the Indiana Chambers. Beni suspects after-pulsing and wants to add methane later. First we will look with iron source (Lubomir) to see if we have after-pulsing, then we will play with the threshold.

- Beni made a code to transform the evio format to hddm. Now Simon can use his powerful code to analyze the test data. Very important step!


- No comments so far received for the FDC section. Eugene: can we find better pictures, drawings; a photo of one complete package will be nice. Lubomir will work on that.